OpenStudy (anonymous):

will medal and fan! Solve for x. 2x-4 ------ = 2x/20 10 --?--

OpenStudy (anonymous):

@harleygirl18 @AlexandervonHumboldt2

OpenStudy (anonymous):

Please help i will take any explanation

OpenStudy (harleygirl18):

20(2x−4)=10(2x) 40x−80=20x

OpenStudy (anonymous):

20x tysm @harleygirl18 !

OpenStudy (anonymous):

20x is correct right? @harleygirl18

OpenStudy (harleygirl18):

yep

OpenStudy (anonymous):

ok cool